Now, three masses m1=3.4kg, m2=10.2kg and m3=6.8kg x=14 hang from three identical springs in a motionless elevator. Finally the elevator is moving downward with a velocity of v=-2.8m/s and also accelerating downward at an acceleration of a=-3.3m/s^2.

The elevator is
A. speeding up
B. slowing down
C. moving at a constant speed

Rank the distances the springs are extended from their unstretched lengths:
A. x1=x2=X3
B.X1>x2>x3
C. x1<x2<x3

what is the distance the Middle spring is extended from its unstretched length?
Answer #1

1) option A speeding up is correct.

it is because, is going downward and velocity is displayed negative which means upward is positive. Similarly acceleration is displayed taking upward direction as positive. Velocity and acceleration have same direction, so speeding up.

2) The more the mass, the more will be extension in order to balance effective weight m(g+a)

order has to be x1<x3<x3.....please recheck the mass data.

3) x3 = m3(g+a)/k = 6.8*(9.8-3.3)/k........not put the value of k, and you will get the extension in length.